OpenStudy (anonymous):

A car averages 40 miles per hour for the first 6 hours of a trip and averages 60 miles per hour for each additional hour of travel time. If the average speed for the entire trip is 55 miles per hour, how many hours long is the trip? (A) 8 (B) 12 (C) 16 (D) 18 (E) 24

OpenStudy (anonymous):

ok, so you know the two averages, so basically what you want to find is the weight of the second average speed. Think of it like calculating your final mark: exam*0.5 + test*.5 = final mark. So you have to do the same thing with the speeds keeping in mind that 55 will be the right side of the equation: 40*6 + 60*unknown = 55. Now you must solve for this unknown.

OpenStudy (anonymous):

240+60x=55 I ended up with a negative..O_O

OpenStudy (anonymous):

oh i forgot to mention, since its an average, you need to divide by the total, so the left side of the equation should be: (40*6 + 60*unknown)/(6+y)=55

OpenStudy (anonymous):

y = unknown

OpenStudy (anonymous):

240+60y -------- =55 6+y Correct so far?

OpenStudy (anonymous):

yep keep solving

OpenStudy (anonymous):

[(40*6)+(60*x)] / [6+x] = 55 240+60x = 55x+330 5x = 90 x = 18